# Agentic Jigsaw Interaction Learning for Enhancing Visual Perception and Reasoning in Vision-Language Models

### v1

@@ 4.2 Main Results
To further assess whether jigsaw training enhances performance on general vision downstream tasks, we evaluate the model on 9 benchmarks: high-resolution image understanding (HRBench4K (Wang et al., 2025c), HRBench8K (Wang et al., 2025c), VStarBench (Wu & Xie, 2024)), real-world scene understanding (MME-RealWorld (Zhang et al., 2024), RealWorldQA (xAI, 2024)), fine-grained visual recognition (MMVP (Tong et al., 2024), BLINK (Fu et al., 2024)), visual reasoning (MMMU (Yue et al., 2024)), and hallucination evaluation (HallusionBench (Guan et al., 2024)). As shown in Table 3, the jigsaw-trained model demonstrates strong visual generalization, achieving notable gains on HRBench4K (+4.2%), HRBench8K (+5.2%), and VStarBench (+4.2%). On average, it surpasses the base model Qwen2.5-VL-7B by 3.1% across all 9 benchmarks, providing compelling evidence that jigsaw-based training effectively enhances the model’s ability to capture visual relations and strengthen reasoning skills, thereby improving its performance on general vision downstream tasks.
@@ Appendix A Data Distribution
$\bullet$ High-resolution and visual search images (39.7%, 6.2K samples): To enhance the model’s fine-grained perception in high-resolution images, we collect data from the VStar (Wu & Xie, 2024) dataset, high-resolution natural scenes in DeepEyes (Zheng et al., 2025), and selected samples from HRBench (Wang et al., 2025c). This improves the model’s ability to capture subtle visual cues and recognize small object attributes.

### v2

@@ 4.2 Main Results
To further assess whether jigsaw training enhances performance on general vision downstream tasks, we evaluate the model on 9 benchmarks: high-resolution image understanding (HRBench4K (Wang et al., 2025c), HRBench8K (Wang et al., 2025c), VStarBench (Wu and Xie, 2024)), real-world scene understanding (MME-RealWorld (Zhang et al., 2024), RealWorldQA (xAI, 2024)), fine-grained visual recognition (MMVP (Tong et al., 2024), BLINK (Fu et al., 2024)), visual reasoning (MMMU (Yue et al., 2024)), and hallucination evaluation (HallusionBench (Guan et al., 2024)). As shown in Table 3, the jigsaw-trained model demonstrates strong visual generalization, achieving notable gains on HRBench4K (+4.2%), HRBench8K (+5.2%), and VStarBench (+4.2%). On average, it surpasses the base model Qwen2.5-VL-7B by 3.1% across all 9 benchmarks, providing compelling evidence that jigsaw-based training effectively enhances the model’s ability to capture visual relations and strengthen reasoning skills, thereby improving its performance on general vision downstream tasks.
@@ Appendix A Data Distribution
$\bullet$ High-resolution and visual search images (39.7%, 6.2K samples): To enhance the model’s fine-grained perception in high-resolution images, we collect data from the VStar (Wu and Xie, 2024) dataset, high-resolution natural scenes in DeepEyes (Zheng et al., 2025), and selected samples from HRBench (Wang et al., 2025c). This improves the model’s ability to capture subtle visual cues and recognize small object attributes.